OpenStudy (anonymous):

Find the average rate of change for the function over the given interval y=4x^2 between x=0 and x=7/4 please explain steps

OpenStudy (anonymous):

your job is to compute \[\frac{f(\frac{7}{4})-f(0)}{\frac{7}{4}-0}\]

OpenStudy (anonymous):

not hard in this example because\(f(0)=0\) giving \[\frac{f(\frac{7}{4})}{\frac{7}{4}}\]

OpenStudy (saifoo.khan):

@satellite73 , it says "rate" so dont we have to find it's derivative?

OpenStudy (anonymous):

\[f(\frac{7}{4})\times \frac{7}{7}=4(\frac{7}{4})^2\times \frac{4}{7}\]

OpenStudy (anonymous):

"average rate"

OpenStudy (anonymous):

so lost

OpenStudy (saifoo.khan):

so rate = derivative? slope? idk..

OpenStudy (anonymous):

\[4\times \frac{7}{4}\times \frac{7}{4}\times \frac{4}{7}=4\times \frac{7}{4}=7\]

OpenStudy (anonymous):

@yomamabf average rate is like the slope of a line

OpenStudy (anonymous):

\[m=\frac{y_2-y_1}{x_2-x_1}\]

OpenStudy (anonymous):

your function is not a line, but you have two points on the curve, namely \((0, f(0))\) and \((\frac{7}{4},f(\frac{7}{4}))\)

OpenStudy (anonymous):

so the answer is 1/3

OpenStudy (anonymous):

since \(f(0)=0\) your two points are \((0,0)\) and \((\frac{7}{4},\frac{49}{4})\)

OpenStudy (anonymous):

no not 1/3, rather \[\frac{\frac{49}{4}}{\frac{7}{4}}=\frac{49}{4}\times \frac{4}{7}=7\]

OpenStudy (anonymous):

thanku